OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

its either the third or fourth one

OpenStudy (aroub):

y=mx+b , m=-6 , (-2,-3) -3=-6(-2)+b -3=12+b -15=b y=-6x-15 so its the first one! :)